# A Counterexample To Hildebrand’s Conjecture On Stable Sets

### v1

@@ Abstract.
We provide a counterexample to a conjecture of Hildebrand which states that if $\mathcal{S}$ has positive lower density and for all $d$, $n$ is in $\mathcal{S}$ if and only if $dn$ is in $\mathcal{S}$ then $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)$ has positive lower density and in particular is nonempty. Heuristically, we define $\mathcal{S}$ by $3^{\ell}n\in\mathcal{S}$ if and only if $\text{Re}(n^{iT}\chi(n))>0$ where $\chi$ is the nontrivial multiplicative character mod 3 and $T$ is chosen such that $p^{iT}\approx\chi(p)$.
@@ 1. Introduction
We say a set $\mathcal{S}$ is stable if for all $p$, for almost every $n$ is the sense of density $n$ is in $\mathcal{S}$ if and only if $pn$ is in $\mathcal{S}$ i.e.
(1) $d(\{n\in\mathbb{N}\colon\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}\}|)=0$
which is to say
$\lim_{N\rightarrow\infty}\frac{|\{n\leq N\colon\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}\}|}{N}=0.$
In general we write $\underline{d}(A)$ for the lower density of $A$ and $\overline{d}(A)$ for the upper density of $A$ i.e.
$\underline{d}(A)=$ $\liminf_{N\rightarrow\infty}\frac{|A\cap[N]|}{N}$
and
$\overline{d}(A)=$ $\limsup_{N\rightarrow\infty}\frac{|A\cap[N]|}{N}$
where $[N]=\{1,2,\ldots,N\}$. We remark that knowing (1) for primes $p$ implies the same for all natural numbers. Hildebrand proved the following conjecture of Balog
@@
$\underline{d}(\mathcal{S}\cap\mathcal{S}+1)>0.$
Applying the theorem to the set $n$ such that the largest prime factor of $n$ is at least $n^{1-\varepsilon}$ Hildebrand ([Hildebrand1]) was able to show that for a positive lower density set  [^1] In fact, the logarithmic density of this set can be computed using Theorem 1.4 of [Joni] of $n$, $n$ and $n+1$ both have prime factors of size at least $n^{1-\varepsilon}$. Hildebrand conjectured
@@
$\underline{d}(\mathcal{S}\cap\mathcal{S}+1\cap\mathcal{S}+2\cap\cdots\cap\mathcal{S}+(k-1))>0.$
@@
$\mathcal{S}\cap\mathcal{S}+1\cap\mathcal{S}+2\cap\cdots\cap\mathcal{S}+(k-1)\neq\emptyset.$
@@
Note that this is equivalent to the conjecture that the upper density of $\mathcal{S}\cap\mathcal{S}+1\cap\mathcal{S}+2\cap\cdots\cap\mathcal{S}+(k-1)$ is positive.Otherwise, given a set where the density is 00, we could delete the set $\mathcal{S}\cap\cdots\cap(\mathcal{S}+k)$ from $\mathcal{S}$ and since that would only effect a set of density 00 it would not change whether $\mathcal{S}$ was a stable set.
Hildebrand ([Hildebrand2]) showed that Conjecture 1.2 holds if $\underline{d}(\mathcal{S})>\frac{k-2}{k-1}$. Tao and Teräväinen ([TJ]) showed a number of variants of Conjecture 1.2 hold. For example if $\mathcal{S}$ is uniformly distributed in small intervals and $\underline{d}(\mathcal{S})>1-\frac{1}{k-\frac{4}{3}+o(1)}$ then the conjecture holds. In this note we prove
@@
Conjecture 1.3 fails for $k=3$ for a set of lower density $\frac{1}{4}$.
We will do this by constructing a set $\mathcal{S}$ for which the intersection has density zero. Applying Remark 1.4, one can obtain a set where the intersection is actually empty.
@@
We use $\underline{d}$, $\overline{d}$ and $d$ for lower density, upper density and density respectively. We say $x=y+O(z)$ if there is a constant $C$ such that $|x-y|\leq C|z|$. We say $x=o_{k\rightarrow\infty}(y)$ if $\lim_{k\rightarrow\infty}\frac{x}{y}=0$. We say $x\lesssim y$ if there exists a constant $C$ such that $x\leq Cy$. We use $[N]$ for $\{1,2,\ldots,N\}$ and $[a,b]$, $(a,b]$ and $(a,b)$ for closed, half-open and open intervals. We set $e(t)=e^{2\pi it}$. Let $\mathbbm{1}_{x\in A}$ denote the indicator function which is $1$ if $x$ is in $A$ and 00 otherwise.
@@ Acknowledgments
I would like to greatly thank Terence Tao for suggesting this problem and sharing his thoughts, in particular that near counterexamples from his work with Matomaki and Radziwill might be useful (see [MRT]). I would also like to thank Joni Teravainen and Nikos Frantzikinakis for their comments on earlier drafts.
@@ 2. Proof of the Main Theorem
$\begin{cases}p^{iT}\approx 1&\text{ whenever }p=1\text{ mod }3\\ p^{iT}\approx-1&\text{ whenever }p=-1\text{ mod }3\end{cases}$
@@
(2) $\begin{cases}3^{\ell}\cdot n\in\mathcal{S}&\text{ whenever }n=1\text{ mod }3\text{ and }\text{Re}(n^{iT})>0\\ 3^{\ell}\cdot n\in\mathcal{S}&\text{ whenever }n=-1\text{ mod }3\text{ and }\text{Re}(n^{iT})<0\\ 3^{\ell}\in\mathcal{S}&\text{ for all }\ell.\end{cases}$
@@
We remark that for fixed $T_{i}$ the error term cannot go to 00 as $N\rightarrow\infty$. For example, suppose we $T_{1}=1$, $T_{2}=\frac{1}{2\log 2}$, $m=2$ and $a_{2}=0$ so that $(2n)^{iT_{2}}=-n^{iT_{2}}$. Then for some choices of $N$ we will have the following points contained in our interval:
@@
which by the geometric series formula should have size $\approx\frac{2}{3}N$. Similarly, for some choices of $N$ the compliment will be our interval which will have size approximately $\frac{1}{3}N$.
@@
Let $T_{1}$ and $T_{2}$ be real numbers greater than $1$ and let $N\geq 2$ and $m$ be natural numbers. Let $I_{1}$ and $I_{2}$ be subsets of $[01]$. Then
|$\frac{|\{n\leq N\colon n^{iT_{1}}\in e(|I_{1}|)\text{ and }n^{iT_{2}}\in e(I_{2})\}|}{N}$
@@
Now we can finally return to the proof of Theorem 1.5. Let $D_{1},D_{2},D_{3},\ldots$ be an increasing sequence of natural numbers and $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\ldots$ be a decreasing sequence of positive real numbers approaching 00. Let $J_{k}=\lceil\frac{1}{\varepsilon}\rceil+1$. By Lemma 2.1, for all real numbers $T_{0}$, there exists $T_{k}\geq T_{0}$ such that for all $p\leq D_{k}$,
(3) $\begin{cases}|p^{iT_{k}}-1|\leq\varepsilon_{k}&\text{ whenever }p=1\text{ mod }3\\ |p^{iT_{k}}+1|\leq\varepsilon_{k}&\text{ whenever }p=-1\text{ mod }3.\end{cases}$
We will fix a sequence of real numbers $T_{k}$ and a sequence of natural numbers $N_{k,0},\ldots,N_{k,2J_{k}}$ such that for each $k$, $T_{k}$ is sufficiently large depending on $\varepsilon_{k}$ and $D_{k}$ and $N_{k,j}$ is sufficiently large depending on $T_{k}$, $T_{k+1}$ and $N_{r,s}$ for $r<k$ and any $s$ or for $r=k$ and any $s<j$. Thus
$D_{k},\frac{1}{\varepsilon_{k}}\ll T_{k}\ll T_{k+1}\ll N_{k,0}\ll N_{k,1}\ll\cdots\ll N_{k,2J_{k}}\ll N_{k+1,1}.$
For convenience define $N_{k}=N_{k,0}$, $N_{k,2J_{k}+1}=N_{k+1}$ and $N_{0}=1$. The precise nature of how big each variable needs to be depending on the last could be worked out by the interested reader but for instance
@@
    $\frac{1}{T_{k}}+\frac{T_{k}\log T_{k}}{T_{k}}+\frac{T_{k}\log T_{k+1}}{N_{k}}$ tends to 00 as $k$ tends to infinity.
@@
    $d(\{n\colon|n-N_{k,j}|<3\text{ for some }k,j\})$
Let $n$ be a natural number in the interval $[N_{k,j},N_{k,j+1})$ or in the interval $[N_{k,j},N_{k+1,0})$ if $j=2J_{k}$ and suppose $n$ is not divisible by $3$. Let $\alpha(n)=0$ for $j\leq J_{k}$ and $\alpha(n)=(j-J_{k})\varepsilon_{k}$ for $j>J_{k}$. Let $\beta(n)=J_{k}-j$ for $j\leq J_{k}$ and $\beta(n)=0$ for $j>J_{k}$. Then we say $3^{\ell}n$ is in $\mathcal{S}$ for all $\ell$ if and only if
$\begin{cases}3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=1\text{ mod }3\text{ and }\text{Re}(n^{iT_{k}})>-\alpha(n)\text{ and }\text{Re}(n^{iT_{k+1}})>-\beta(n)\\ 3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=-1\text{ mod }3\text{ and}\text{Re}(n^{iT_{k}})<\alpha(n)\text{ and }\text{Re}(n^{iT_{k+1}})<\beta(n)\\ \end{cases}$
@@
When $\alpha>1$ then the conditions depending on $\alpha$ are trivially satisfied. Similarly for the conditions depending on $\beta$ when $\beta>1$. Thus, by slowly changing $\alpha$ and $\beta$ we slowly “turn on” and “turn off” the restrictions that $\text{Re}(n^{iT_{k}})>0$ and $\text{Re}(n^{iT_{k+1}})>0$.
@@
Let $p$ be a prime. If $p=3$ then by examining the definition of $\mathcal{S}$, $n$ is in $\mathcal{S}$ if and only if $3n$ is in $\mathcal{S}$ so assume $n\neq 3$. Let $k$ be such that $p\leq D_{k}$ which exists since $D_{k}$ tends to infinity by assumption. Consider the case of some $n$ in $[N_{k,j},N_{k,j+1})$ which is not divisible by $3$. Then if $j<2J$, by (2) either $pn$ is in $[N_{k,j},N_{k,j+1})$ or in $[N_{k,j+1},N_{k,j+2})$. In this case
$\begin{cases}3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=1\text{ mod }3\text{ and }\text{Re}(n^{iT_{k}})>-\alpha(n)\text{ and }\text{Re}(n^{iT_{k+1}})>-\beta(n)\\ 3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=-1\text{ mod }3\text{ and}\text{Re}(n^{iT_{k}})<\alpha(n)\text{ and }\text{Re}(n^{iT_{k+1}})<\beta(n)\\ \end{cases}$
and
$\begin{cases}3^{\ell}\cdot pn\in\mathcal{S}_{k}&\text{ whenever }pn=1\text{ mod }3\text{ and }\text{Re}((pn)^{iT_{k}})>-\alpha(pn)\text{ and }\text{Re}((pn)^{iT_{k+1}})>-\beta(pn)\\ 3^{\ell}\cdot pn\in\mathcal{S}_{k}&\text{ whenever }pn=-1\text{ mod }3\text{ and }\text{Re}((pn)^{iT_{k}})<\alpha(pn)\text{ and }\text{Re}((pn)^{iT_{k+1}})<\beta(pn).\\ \end{cases}$
Importantly, $\alpha$ and $\beta$ have not changed very much – each has changed by at most $\varepsilon_{k}$ by definition of $\alpha$ and $\beta$. If $p=1\text{ mod }3$ then
$\text{Re}((pn)^{iT_{k}})-\alpha(pn)=$ $\text{Re}(n^{iT_{k}})+O(\varepsilon_{k})-\alpha(pn)$
$=$ $\text{Re}(n^{iT_{k}})+O(\varepsilon_{k})-\alpha(n)$
where the first step is by (3) and the second step is because $\alpha$ changes slowly. Thus, $\text{Re}((pn)^{iT_{k}})-\alpha(pn)$ is positive if and only if $\text{Re}(n^{iT_{k}})-\alpha(n)$ unless $n^{iT_{k}}$ lies in one of two small intervals corresponding to the $\varepsilon_{k}$ neighborhood of the two points for which $\text{Re}(n^{iT_{k}})=\alpha(n)$. Similarly, the sign of $\text{Re}((pn)^{iT_{k+1}})-\beta(pn)$ is the same as the sign of $\text{Re}(n^{iT_{k+1}})-\beta(n)$ unless $n$ satisfies $\text{Re}(n^{iT_{k+1}})=\beta(n)+O(\varepsilon_{k})$. By Lemma 2.1 we can estimate the probability that either of these happen. In particular by (1) this probability goes to 00. The other important case to consider is what happens if $n$ is in $[N_{k,2J_{k}},N_{k+1,0})$ and $pn$ is in $[N_{k+1,0},N_{k+1,1})$. In this case, notice that the $\alpha$-condition for $n$ and the $\beta$-condition for $pn$ actually play no role because each is of the form $\text{Re}(z)\geq-1$ or $\text{Re}(z)\leq 1$ which are trivially true. Thus, in both cases
$\begin{cases}3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=1\text{ mod }3\text{ and }\text{Re}(n^{iT_{k+1}})>0\\ 3^{\ell}\cdot n\in\mathcal{S}_{k}&\text{ whenever }n=-1\text{ mod }3\text{ and}\text{Re}(n^{iT_{k+1}})<0\\ \end{cases}$
and
$\begin{cases}3^{\ell}\cdot pn\in\mathcal{S}_{k}&\text{ whenever }pn=1\text{ mod }3\text{ and }\text{Re}((pn)^{iT_{k+1}})>0\\ 3^{\ell}\cdot pn\in\mathcal{S}_{k}&\text{ whenever }pn=-1\text{ mod }3\text{ and }\text{Re}((pn)^{iT_{k+1}})<0.\\ \end{cases}$
[…]
Now we only handled the case of $n$ not divisible by $3$. Let’s see how this allows us to estimate the number of $n$ such that $\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}$. Let $\ell$ be a natural number. Let $E(A,B)$ denote the number of $n$ between $A$ and $B$ such that $\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}$ and let $E_{0}(A,B)$ be the number for which $n$ is not divisible by $3$. Then we can write $E(1,N)$ as a kind of Cesàro average:
@@
since each $n\geq N3^{-j}$ can show up $j$ times until $3^{j}n\geq N$ i.e. we have the expression $\mathbbm{1}_{3^{h}n\in\mathcal{S}}\neq\mathbbm{1}_{p3^{h}n\in\mathcal{S}}$ for each $h\leq j$. The remaining $O\left(\frac{N}{3^{\ell}}\right)$ accounts for all numbers divisible by $3^{\ell}$ and therefore not included in our count and also for all the numbers between $1$ and $\frac{N}{3^{\ell}}$. We conclude that if the frequency of $\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}$ in $[3^{-j}N,3^{-j+1}N)$ for $j\leq\ell$ and for $n$ not divisible by $3$ is at most $\varepsilon_{k}$ then at a minimum

### v2

@@ Abstract.
We provide a counterexample to a conjecture of Hildebrand which states that if $\mathcal{S}$ has positive lower density and is stable i.e. for all $d$, $n$ is in $\mathcal{S}$ if and only if $dn$ is in $\mathcal{S}$ except on a set of density 0 then $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)$ has positive lower density and in particular is nonempty. We further show there exists a stable set of density $1-\frac{1}{q-1}$ such that $\mathcal{S}\cap\cdots\cap(\mathcal{S}+q-1)=\emptyset$ when $q$ is a prime matching a bound proven by Hildebrand. Finally, we construct a function $f\colon\mathbb{N}\rightarrow\{\pm 1\}$ such that f(pn)=f(n)f(pn)=-f(n) for all but a 0 density set of $n$ depending on the prime $p$ but which fails the analogues of Sarnak and Chowla’s conjectures.
@@ 1. Introduction
Erdős ([Erdos1976]) asked whether there are infinitely many twin smooth numbers, numbers $n$ and $n+1$ all of whose prime factors are less than $n^{\varepsilon}$ for some fixed $\varepsilon>0$. Hildebrand ([Hildebrand1]) solved the question, showing not only that there are infinitely many twin smooth numbers, but that the set of such numbers has positive lower density. In fact Hildebrand showed something much more general.
We say a set $\mathcal{S}$ is stable if for all $p$, for all but a zero density set of $n$, $n$ is in $\mathcal{S}$ if and only if $pn$ is in $\mathcal{S}$ i.e.
(1) $d(\{n\in\mathbb{N}\colon\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}\})=0$
We remark that knowing (1) for primes $p$ implies the same for all natural numbers. We also remark that examples of stables sets include:
  1. (1)
    $n^{\varepsilon}$-smooth numbers $n$ for any fixed value of $\varepsilon$
  2. (2)
    any set of density 0
  3. (3)
    the set of numbers $n$ for which $\frac{\Omega(n)-\log\log n}{\sqrt{\log\log n}}$ is between 100 and 101
  4. (4)
    the set of numbers $n$ which have more prime factors which are $1\text{ mod }10$ than $9\text{ mod }10$ and more prime factors which are $1\text{ mod }3$ than $2\text{ mod }3$
  5. (5)
    the intersection of any finite number of stable sets.
Hildebrand proved the following remarkable conjecture of Balog:
@@
$\underline{d}(\mathcal{S}\cap(\mathcal{S}+1))>0.$
It is perhaps worth remarking that [MR1619809] later found a specific infinite sequence of $k$ consecutive smooth numbers for any $k$ and [Joni] show that with logarithmic density the probability that $n$ and $n+1$ are smooth is the product of the probabilities. Hildebrand ([Hildebrand2]) was also able to show to following generalization of Theorem 1.1
Theorem 1.2 ([Hildebrand2]).
Let $k$ be a natural number and let $\mathcal{S}$ be a stable set with
$\underline{d}(\mathcal{S})>1-\frac{1}{k-1}.$
Then for any integers $h_{1},\ldots,h_{k}$
$\underline{d}((\mathcal{S}+h_{1})\cap\cdots\cap(\mathcal{S}+h_{k}))>0.$
This led Hildebrand to conjecture
@@
$\underline{d}(\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)\cap\cdots\cap(\mathcal{S}+(k-1)))>0.$
@@
$\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)\cap\cdots\cap(\mathcal{S}+(k-1))\neq\emptyset.$
@@
Note that this is equivalent to the conjecture that the upper density of $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)\cap\cdots\cap(\mathcal{S}+(k-1))$ is positive. Otherwise, given a set where the density is 0, we could delete the set $\mathcal{S}\cap\cdots\cap(\mathcal{S}+k)$ from $\mathcal{S}$ and since that would only affect a set of density 0 it would not change whether $\mathcal{S}$ was a stable set.
Tao and Teräväinen ([TJ]) showed a number of variants of Conjecture 1.3 hold. For example if $\mathcal{S}$ is uniformly distributed in small intervals and $\underline{d}(\mathcal{S})>1-\frac{1}{k-\frac{4}{3}+o(1)}$ then the conjecture holds. In this note we prove
@@
Conjecture 1.4 fails for $k=3$ for a set of density $\frac{1}{2}$.
We will do this by constructing a set $\mathcal{S}$ for which the intersection has density zero. Applying Remark 1.5, one can obtain a set where the intersection is actually empty. More generally, we will show
Theorem 1.7.
Conjecture 1.4 fails for any prime $k$ for a set of density $1-\frac{1}{k-1}$.
We remark that this perfectly matches Hildebrand’s bound, meaning this is essentially best possible.
We also give a construction which may be of interest to those studying multiplicative functions. Note that a stable set can be thought of as a binary valued function $f$ satisfying $f(pn)=f(n)$ for all but a density 0 set of $n$ depending on $p$. We can likewise use a similar construction to prove the following.
Theorem 1.8.
There exists a function $f\colon\mathbb{N}\rightarrow\{\pm 1\}$ such that, for any prime $p$, the set
$\{n\in\mathbb{N}\colon f(n)\neq-f(pn)\}$
has density 0 and for which
$\frac{1}{x}\sum_{n\leq x}f(n)f(n+1)<-\varepsilon<0$
for some $\varepsilon$ and for all $x$ sufficiently large. Furthermore
$\lim_{H\rightarrow\infty}\lim_{x\rightarrow\infty}\frac{1}{x}\sum_{n\leq x}\left|\frac{1}{H}\sum_{h\leq H}f(n)\mathbbm{1}_{n=1\text{ mod }3}\right|=\frac{1}{3}.$
We remark that the construction is somewhat similar to constructions in [MR3435814] and [MRT]. The property that f(pn)=f(n)f(pn)=-f(n) except on a set of density 0 means $f$ behaves somewhat like the Liouville function, the completely multiplicative function $\lambda$ such that $\lambda(pn)=-\lambda(n)$ for any prime $p$ and any $n$. The Chowla and Sarnak conjectures for the Liouville function suggest that $\lambda$ does not correlate with any of its shifts nor with any entropy 0 sequence. However, Theorem 1.8 suggests that there exists a function $f$ such that f(pn)=f(n)f(pn)=-f(n) 100% of the time but which fails the Chowla and Sarnak conjectures. Thus, it is not enough to use that f(pn)=f(n)f(pn)=-f(n) 100% of the time in order to prove those conjectures. In particular, there is a method introduced to the field by Bourgain, Sarnak and Ziegler ([BSZ]) sometimes called the Katai criterion, the Turan-Kubilius method, or some combination of the six names which has been used to show many instances of the Sarnak conjecture. And yet, the Bourgain-Sarnak-Ziegler-Katai-Turan-Kubilius method works for any function which satisfies f(pn)=f(n)f(pn)=-f(n) 100% of the time. Therefore, Theorem 1.8 rules out this method by itself and others like it when it comes to proving the full Sarnak or Chowla conjectures.
@@
We use $\underline{d}$, $\overline{d}$ and $d$ for lower density, upper density and density respectively. We say $x=y+O(z)$ if there is a constant $C$ such that $|x-y|\leq C|z|$. We say $x=o_{k\rightarrow\infty}(y)$ if $\lim_{k\rightarrow\infty}\frac{x}{y}=0$. We say $x\lesssim y$ if there exists a constant $C$ such that $x\leq Cy$. We use $[N]$ for $\{1,2,\ldots,N\}$ and $[a,b]$, $(a,b]$ and $(a,b)$ for closed, half-open and open intervals. We set $e(t)=e^{2\pi it}$. Let $\mathbbm{1}_{x\in A}$ denote the indicator function which is $1$ if $x$ is in $A$ and 0 otherwise.
@@ Acknowledgments
I would like to greatly thank Terence Tao for suggesting this problem and sharing his thoughts, in particular that near counterexamples from his work with Matomaki and Radziwill might be useful (see [MRT]). I would also like to thank Joni Teräväinen and Nikos Frantzikinakis for their comments on earlier drafts. Supported by the research grant ELIDEK-HFRI-NextGenerationEU-15689.
@@ 2. Proof of the Main Theorem
[…]
@@
Since $n^{iT}\approx(n+1)^{iT}$ for large $n$, for most $n$ either $\text{Re}(n^{iT})>0,\text{Re}((n+1)^{iT})>0$ and $\text{Re}((n+2)^{iT})>0$, one of which is $1\text{ mod }3$ and therefore not in $\mathcal{S}$ or $\text{Re}(n^{iT})<0,\text{Re}((n+1)^{iT})<0$ and $\text{Re}((n+2)^{iT})<0$, one of which is $1\text{ mod }3$ and therefore not in $\mathcal{S}$. Thus, $\mathcal{S}\cap(\mathcal{S}+1)\cap(\mathcal{S}+2)$ will have density 0. Furthermore, for any $p=1\text{ mod }3$ we have $(pn)^{iT}=p^{iT}n^{iT}\approx n^{iT}$ by choice of $T$, so for most $n$ multiplying by $p$ does not change whether $n$ is $1$ or $1\text{ mod }3$ and also does not change whether $\text{Re}(n^{iT})>0$. Thus, $n$ is in $\mathcal{S}$ if and only if $pn$ is in $\mathcal{S}$ unless $\text{Re}(n)\approx 0$. Similarly for $p=-1\text{ mod }3$, $\text{Re}((pn)^{iT})\approx-\text{Re}(n^{iT})$ so for most $n$ not divisible by $3$ multiplying by $p$ swaps whether $n$ is $1$ or $1\text{ mod }3$ and also swaps whether $\text{Re}(n^{iT})>0$ so again $n$ is in $\mathcal{S}$ if and only if $pn$ is in $\mathcal{S}$.
@@
We remark that for fixed $T_{i}$ the error term cannot go to 0 as $N\rightarrow\infty$. For example, suppose we $T_{1}=1$, $T_{2}=\frac{1}{2\log 2}$, $m=2$ and $a_{2}=0$ so that $(2n)^{iT_{2}}=-n^{iT_{2}}$. Then for some choices of $N$ we will have the following points contained in our interval:
@@
which by the geometric series formula should have size $\approx\frac{2}{3}N$. Similarly, for some choices of $N$ the complement will be our interval which will have size approximately $\frac{1}{3}N$.
@@
Let $T_{1}$ and $T_{2}$ be real numbers greater than $1$ and let $N\geq 2$ and $m$ be natural numbers. Let $I_{1}$ and $I_{2}$ be subsets of $[0,1]$. Then
$\frac{|\{n\leq N\colon n^{iT_{1}}\in e(I_{1})\text{ and }n^{iT_{2}}\in e(I_{2})\}|}{N}$
@@
Now we can finally return to the proof of Theorem 1.6. Let $D_{1},D_{2},D_{3},\ldots$ be an increasing sequence of natural numbers and $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\ldots$ be a decreasing sequence of positive real numbers approaching 0. Let $J_{k}=\lceil\frac{2}{\varepsilon_{k}}\rceil+1$. By Lemma 2.1, for all real numbers $T_{0}$, there exists $T_{k}\geq T_{0}$ such that for all $p\leq D_{k}$,
[…]
We can also fix $t_{k}\geq T_{k}$ such that
(4) $|p^{it_{k}}-1|\leq\varepsilon_{k}$
for all $p\leq D_{k}$.
We will fix two sequences of real numbers $T_{k}$ and $t_{k}$ and a sequence of natural numbers $N_{k,0},\ldots,N_{k,J_{k}}$ such that for each $k$, $T_{k}$ is sufficiently large depending on $\varepsilon_{k}$ and $D_{k}$, $t_{k}$ is sufficiently large depending on $T_{k}$, $D_{k}$ and $\varepsilon_{k}$ and $N_{k,j}$ is sufficiently large depending on $T_{k}$, $T_{k+1}$, $t_{k}$, $\varepsilon_{k}$ and $N_{r,s}$ for $r<k$ and any $s$ or for $r=k$ and any $s<j$. Thus
$D_{k},\frac{1}{\varepsilon_{k}}\ll T_{k}\ll t_{k}\ll T_{k+1}\ll N_{k,0}\ll N_{k,1}\ll\cdots\ll N_{k,J_{k}}\ll N_{k+1,1}.$
For convenience define $N_{k}=N_{k,0}$, $N_{k,J_{k}+1}=N_{k+1}$ and $N_{0}=1$. The precise nature of how big each variable needs to be depending on the last could be worked out by the interested reader but for instance
@@
    $\frac{1}{T_{k}}+\frac{T_{k}\log T_{k}}{t_{k}}+\frac{T_{k}\log T_{k+1}}{N_{k}}$ tends to 0 as $k$ tends to infinity.
@@
    $d(\{n\colon|n-N_{k,j}|<3\text{ for some }k,j\})=0.$
Let $n$ be a natural number in the interval $N_{k,j},N_{k,j+1})$ or in the interval $N_{k,j},N_{k+1,0})$ if $j=J_{k}$ and suppose $n$ is not divisible by $3$. Let $\alpha(n)=(\varepsilon_{k}\cdot j)-1$. Let
[…]
Then we say $3^{\ell}n$ is in $\mathcal{S}$ for all $\ell$ if and only if
[…]
@@
Whether you use the condition $\chi(n)\cdot\text{ sign}(\text{Re}(n^{iT_{k}}))>0$ or $\chi(n)\cdot\text{ sign}(\text{Re}(n^{iT_{k+1}}))>0$ depends whether $\text{Re}(n^{it_{k}})\geq\alpha(n)$. When $n$ is in $N_{k,0},N_{k,1})$ and $\alpha(n)=-1$, then the conditions that $\text{Re}(n^{it_{k}})\geq\alpha(n)$ is trivially satisfied and when $\alpha(n)>1$ it is necessarily false. Thus, by slowly changing $\alpha$ we slowly “turn on” and “turn off” the restrictions that $\chi(n)\text{Re}(n^{iT_{k}})>0$ and $\chi(n)\text{Re}(n^{iT_{k+1}})>0$.
@@
Let $p$ be a prime. If $p=3$ then by examining the definition of $\mathcal{S}$, $n$ is in $\mathcal{S}$ if and only if $3n$ is in $\mathcal{S}$ so assume $p\neq 3$. Let $k$ be such that $p\leq D_{k}$ which exists since $D_{k}$ tends to infinity by assumption. Consider the case of some $n$ in $N_{k,j},N_{k,j+1})$ which is not divisible by $3$. Then if $j<J_{k}$, by (2) either $pn$ is in $N_{k,j},N_{k,j+1})$ or in $N_{k,j+1},N_{k,j+2})$. Observe that, by (3), we have
$\chi(pn)\text{Re}((pn)^{iT_{k}})=$ $\text{Re}(n^{iT_{k}}p^{iT_{k}}\chi(p))$
$=$ $\text{Re}(n^{iT_{k}}(1+O(\varepsilon_{k})))$
$=$ $\text{Re}(n^{iT_{k}})+O(\varepsilon_{k})$
and similarly
$\chi(pn)\text{Re}((pn)^{iT_{k+1}})=$ $\text{Re}(n^{iT_{k+1}}p^{iT_{k+1}}\chi(p))$
$=$ $\text{Re}(n^{iT_{k+1}})+O(\varepsilon_{k})$
and also because $\alpha(pn)=\alpha(n)+O(\varepsilon_{k})$ and (4) we have
$\text{Re}((pn)^{it_{k}})-\alpha(pn)=$ $\text{Re}(n^{it_{k}}p^{it_{k}})-\alpha(n)+O(\varepsilon_{k})$
$=$ $\text{Re}(n^{it_{k}})(1+O(\varepsilon_{k}))-\alpha(n)+O(\varepsilon_{k}).$
Therefore, each of these expressions can only change signs if one of $n^{iT_{k}}$, $n^{iT_{k+1}}$, or $n^{it_{k}}$ is in one of two small intervals of length $O(\varepsilon_{k})$. By Lemma 2.1 we can estimate the probability that any of these happen. In particular by (1) that probability goes to 0.
The other important case to consider is what happens if $n$ is in $N_{k,2J_{k}},N_{k+1,0})$ and $pn$ is in $N_{k+1,0},N_{k+1,1})$. In this case, notice that the $\text{Re}(n^{it_{k}})<\alpha_{k}(n)$ for $n$ and the $\text{Re}((pn)^{it_{k}})\geq\alpha_{k+1}(pn)$ actually play no role because each is of the form $\text{Re}(z)\geq-1$ or $\text{Re}(z)\leq 1$ which are trivially true. Thus as before $n$ is in $\mathcal{S}$ if and only if $pn$ is in $\mathcal{S}$ unless $\text{Re}(n^{iT_{k+1}})=O(\varepsilon_{k})$ which is rare.
So far, we only handled the case of $n$ not divisible by $3$. Let’s see how this allows us to estimate the number of $n$ such that $\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}$ in general. Let $\ell$ be a natural number. Let $E(A,B)$ denote the number of $n$ between $A$ and $B$ such that $\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}$ and let $E_{0}(A,B)$ be the number of such $n$ which are not divisible by $3$. Then we can write $E(1,N)$ as a kind of Cesàro average:
@@
since each $n\geq N3^{-j}$ can show up $j$ times until $3^{j}n\geq N$ i.e. we have the expression $\mathbbm{1}_{3^{h}n\in\mathcal{S}}\neq\mathbbm{1}_{p3^{h}n\in\mathcal{S}}$ for each $h\leq j$. The remaining $O\left(\frac{N}{3^{\ell}}\right)$ accounts for all numbers divisible by $3^{\ell}$ and therefore not included in our count and also for all the numbers between $1$ and $\frac{N}{3^{\ell}}$. We conclude that if the frequency of $\mathbbm{1}_{n\in\mathcal{S}}\neq\mathbbm{1}_{pn\in\mathcal{S}}$ in $3^{-j}N,3^{-j+1}N)$ for $j\leq\ell$ and for $n$ not divisible by $3$ is at most $\varepsilon_{k}$ then at a minimum